OpenStudy (anonymous):

solving limits at infinity(x->infinity), I have a problem x/sqrt(x^2+1). I tried L'Hospital's rule, and derivative, after simplification, ends up being another form of x/x or infinity/infinity. Is this true? or have I missed something?

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{x}{\sqrt{x^2+1}} \cdot \frac{\frac{1}{\sqrt{x^2}}}{\frac{1}{\sqrt{x^2}}}\]

OpenStudy (amistre64):

if so, then lhopital it again

OpenStudy (anonymous):

You can perform the derivative again.

myininaya (myininaya):

ok remember the |x| is x if x>0 |x| is -x if x<0 since x->positive infinity then we have |x|=x

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{\frac{x}{x}}{\sqrt{\frac{x^2+1}{x^2}}}\]

myininaya (myininaya):

\[\lim_{x \rightarrow \infty}\frac{1}{\sqrt{1+\frac{1}{x^2}}}\]

myininaya (myininaya):

\[\frac{1}{\sqrt{1+0}}=\frac{1}{\sqrt{1}}=\frac{1}{1}=1\]

myininaya (myininaya):

thats without lhospital

myininaya (myininaya):

but you use that guy if you want to

OpenStudy (anonymous):

thank you, that was what I was doing wrong. I should have tried to simplify it without L'hospital If you use L'Hospital, because of the sqrt and chain rule, you always end up with continued -1/2 exponents.